Wednesday, August 23, 2017

From HallCWiki
Jump to: navigation, search

Meeting Information

  • Location CC L210A, 11 am
  • DocDB

Blue Jeans Connection Information

  1. Enter Meeting ID: 895800575
  2. Press #

Agenda

  1. Commissioning plan updates from detector systems
  2. Status update from detector owners, calibrations, efficiencies, etc.
    1. Checking Keithley Current Source with a Newly Calibrated Multimeter - Deepak Bhetuwal
    2. Recent Trigger/DAQ Work -Carlos Yero
    3. Q2 Measurements Update - Holly
  3. Review systems owners and update information
  4. Address system specific points of interest for the Fall 2017 run

Minutes

Beam Line

  1. Update on BCM Studies and Calibrations by Deepak

See Link. Checking Keithley Current Source with a Newly Calibrated Multimeter - Deepak Bhetuwal

Trigger/DAQ

  1. Recent update on Trigger/DAQ in preparation for Hall C Fall-2017/Spring-2018

commissioning experiments. Main work has been to clean up unnecessary cable, and re-route some badly placed cables

  1. EDTM electronics for HMS/SHMS hodoscopes has been set-up and verified to be working
  2. Live Times for the HMS/SHMS were very different at EDTM>500 Hz, this might be due to a

update on the HMS DAQ firmware. Steve Wood will pass by the Counting Room to help on this issue. See Link. Recent Trigger/DAQ Work -Carlos Yero

SHMS

Magnets

  1. characterize Q2 behavior : checking for cross-talk polarity asymmetries, linearity
  2. Dave Mack is skeptical about the errors obtained in the measurement of Q2 field using the Hall probe.
  3. The hall probe might not work as it should for such strong magnetic field gradients according to D. Mack

See Link. Q2 Measurements Update - Holly Szumila